Circular No. 8634

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET C/2005 W3 (KOWALSKI) Richard Kowalski reports a comet discovery on Catalina Sky Survey

0.68-m Schmidt images (discovery observation tabulated below). Coma diameters and tail lengths from CCD images: Nov. 25.32 UT, ∼ 20′′, — in p.a. ∼ 280^{o} (Kowalski); 25.4, 10′′ (well-condensed), 20′′ in p.a. 280^{o}– 290^{o} (E. J. Christensen, Mt. Lemmon 1.5-m reflector; four co-added 30-s frames); 26.3, — (diffuse, extended), ∼ 20′′ toward the west-northwest (J. G. Ries, McDonald Observatory 0.76-m reflector); 27.1, 6′′ (concentrated), 14′′ (straight, broad, 8′′ wide) in p.a. 305^{o} (P. Birtwhistle, Great Shefford, U.K., 0.40-m reflector). The available astrometry (including prediscovery NEAT observations from Nov. 24), preliminary parabolic elements (T = 2005 Apr. 22.927 TT, q = 2.68255 AU, ω = 155^{o}.860, Ω = 220^{o}.463, i = 21^{o}.366, equinox 2000.0), and an ephemeris appear on MPEC 2005-W72.

2005 UT α_{2000} δ_{2000} Mag. Observer
Nov. 25.31896 5^{h}02^{m}28.^{s}03 +6^{o}52′30′′.0  18.2 Kowalski

SUPERNOVAE 2005kl–2005ku The discoveries of twelve supernovae have been reported: 2005kl by

M. Migliardi (from unfiltered CROSS CCD images; cf. IAUC 7373, 8588; via A. Dimai, Cortina d’Ampezzo, Italy), 2005km by J. Burket and W. Li (LOSS/KAIT; cf. IAUC 8632), and 2005kn–2005ku by J. Frieman on behalf of the Sloan Digital Sky Survey II collaboration (all found at mag > 20.4, with 2005kt and 2005ku predicted to peak around now at g = 18.2 and 17.5, respectively; details given on CBET 304).

SN  2005 UT  α_{2000}  δ_{2000}  Mag.  Offset
2005kl Nov. 22.14 12^{h}24^{m}35.^{s}68 +39^{o}23′03′′.5 14.6 6′′ W, 4′′ N
2005km Nov. 22.52 11 20 48.72 +32 17 25.2 17.5 3′′.5 E, 1′′.0 N
2005kt Nov. 7  1 10 58.05 + 0 16 34.1 20.5  —
2005ku Nov. 10  22 59 42.61 − 0 00 49.3 20.7  —

Nothing is visible at the location of 2005kl on Palomar Sky Survey infrared, red, and blue plates. Additional approximate magnitudes for 2005kl in NGC 4369: 2003 May 5, [18.5; 2005 Nov. 23.15 UT, 14.7. Additional KAIT magnitudes for 2005km in UGC 6356: Mar. 17.20, [19.5; Nov. 23.54, 17.3. SNe 2005kr and 2005ks are type-Ib/c supernovae with broad lines similar to that seen in 1998bw; the rest appear to be type-Ia supernovae except for type-Ic supernova 2005kl (cf. CBET 305, 306).

2005 November 27  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2005 CBAT  Daniel W. E. Green